Determine f (xo), * (XO) and fi (xo) for the given point xo if y = P(x) is a solution of the given initial value problem.
Answer #1

a yltayty=0 : 400)=3 :Yc0)=5: Absin converting a Inlit : Y(0) = 3: 4101-5 0/6)=0 y++y+ y =0 ce Posao y+y+ 4y = 0 o Co=